# Generative AI-Driven Human Digital Twin in IoT-Healthcare: A Comprehensive Survey

### v1

@@ III-B GAI-enabled Communication
HDT relies on real-time data transmission to keep synchronization between any PT-VT pair to ensure high-fidelity of VT [8]. This synchronization is, however, data-driven and delay-sensitive. Furthermore, data acquired in the physical world is often massive and complex. In addition to real-time synchronization, interacting with VTs involves more complex information that needs to be transmitted between the PT and users in the physical world. Multi-modal information, such as 3D virtual items, text, images, haptic signals, among others, needs to be transmitted in HDT under various applications, such as the virtual surgery, to enhance the immersive experience. These specific characteristics place a significant burden on current communication networks. To address these issues, this subsection delves into the GAI-aid semantic communication [52] and cross-modal communication [54] for communication in HDT, as summarized in Table IV.
TABLE IV: Summary of GAI-based Approach for Communication in HDT.
@@ III-C GAI-enabled Data Management
Data management is a crucial step in the successful implementation of HDT. First, the collected multi-source data in HDT may have the characteristics of heterogeneity, multi-scale and high noises. Hence, as the critical step in data management, data pre-processing is indispensable in HDT, such that issues like missing data and noise data can be properly handled, and thereby providing high-quality data for downstream tasks in HDT [8]. Moreover, the data in HDT are highly sensitive, especially for individual-level data, and any leakage may result in serious ethical and moral concerns. Therefore, effective security and privacy schemes are imperative to protect data in HDT. In the following, we will discuss the application of GAI in data pre-processing, followed by its application in security and privacy schemes, as summarized in Table V.
TABLE V: Summary of GAI-based Approach for Data Management in HDT.
@@
Traditional data imputation methods, such as K-nearest neighbors-based [77] and deep learning-based imputation [78], often rely on existing observed data for modeling and prediction, and cannot directly generate new data samples. This will make it hard for traditional data imputation methods to handle missing severe data issues, such as data modality missing. To this end, GAI, with its powerful generative capabilities, can effectively handle the missing data issues in HDT. For instance, with the potential of substituting missing data accurately and efficiently, Dong et al. in [79] used GAN to impute missing values in large clinical datasets collected from PTs with mixed-type variables. Specifically, the method adopted by authors was generative adversarial imputation nets (GAIN), where the generator observed some components of a real clinical data vector, imputed the missing components conditioned on what was actually observed, and outputs a complete clinical data vector. The discriminator then took a completed clinical data vector and attempted to determine which components were observed and which were imputed. Additionally, to ensure that discriminator forced generator to learn the desired data distribution, the authors provided discriminator with some additional information in the form of a hint vector. The experimental results showed that GAIN outperformed the traditional imputation models, MICE [80] and missForest [81], in terms of accuracy in the imputation of missing data in clinical datasets, particularly for imbalanced and skewed data, and when the missingness rate was high (50%). Acquiring multi-modality data is crucial for the implementation of HDT. However, the missing modalities may happen in some conditions, such as failure in data transmission. Fortunately, diffusion models have shown favorable results for generating missing modalities utilizing cross-modalities and producing ones using other modality types. For instance, Lyu et al. in [82] proposed a GAI framework, called diffusion and score-matching models, which took advantage of the recently introduced denoising diffusion probabilistic models (DDPMs) [83] and score-based diffusion models [84], for translating MRI to CT. Specifically, they presented conditional DDPM and conditional stochastic differential equation (SDE) [85], where their reverse process was conditioned on T2w MRI images. The authors adopted the DDPM and SDE with three different sampling methods, namely Euler-Maruyama, Prediction-Corrector, and probability flow ordinary differential equation. Their extensive experiments on the Gold Atlas male pelvis dataset [86] demonstrated that the proposed diffusion models outperformed both GAN and CNN-based methods [87] regarding structural similarly index measure and peak signal-to-noise ratio. Similarly, to cope with the missing modality issue, Meng et al. in [48] proposed a unified multi-modal conditional score-based generative approach (UMM-CSGM), which synthesized the missing modality based on all remaining modalities as conditions. The proposed model was a conditional SDE format, employing only a score-based network to learn different cross-modal conditional distributions. The experimental results showed that the UMM-CSGM could generate missing-modality images with higher fidelity and structural information of the brain tissue compared to GAN-based methods [88, 89, 90, 91, 92].
[…]
@@
HDT environments may suffer from various security threats, such as eavesdropping attacks in data transmission between PTs and VTs causing privacy leakage. A typical privacy protection mechanism for HDT data is anonymization [98, 99]. However, this method can easily compromise the data distribution, and it is hard to balance data privacy and usability. To this end, GAI can generate synthesized HDT data, which maximally preserves original data distribution while guaranteeing data security and privacy. For instance, to lower the risk of breaching individual confidentiality during data sharing in HDT, Yoon et al. in [100] proposed a GAN-based framework, called anonymization through data synthesis using GAN (ADS-GAN), for generating synthetic data that closely approximates the distribution of variables in an original EHR dataset, achieving anonymization of EHR, as shown in Fig. 6. Specifically, ADS-GAN generated the synthetic data conditioned on the original data, and different from traditional conditional GAN framework, conditioning variables in ADS-GAN were not pre-determined but instead were optimized from real EHR. Therefore, the ADS-GAN can generate better quality synthetic data than the traditional conditional GAN frameworks [87, 101, 102, 103] while minimizing individual identifiability. However, such methods have limitations regarding the fundamental aspects of real-world EHR data, such as dealing with missing features, varying feature length, categorical features and static features (beyond time series). These fundamental challenges require a holistic re-design in GAN-based synthetic data generation systems. To this end, Yoon et al. in [104] further designed, EHR-Safe, that can jointly represent these diverse data modalities while preserving the privacy of EHR data. Specifically, EHR-Safe was based on a two-stage model that consisted of sequential encoder-decoder networks and GANs. To circumvent the heterogeneity of EHR data mentioned above, the authors used the sequential encoder-decoder networks to learn the mapping from the raw EHR data to low-dimensional representations and vice versa, and the mapped information is used for GAN training. The experimental results showed that the fidelity of synthetic data generated by EHR-Safe is almost-identical with real data, while yielding ideal performance in practical privacy metrics.
@@ III-D GAI-enabled Digital Modeling
The classic HDT digital modeling technology, mechanistic modeling, has pioneered the integration of biology and physiology domain knowledge to allow robust and accurate modeling [105]. For instance, a human heart DT model has been developed, which consisting of approximately 100 million virtual heart cell patches, with each patch modeled by around 50 equations [106]. The human heart DT accurately represented the interconnected cardiac muscle cells, effectively simulating the transmission of electrical currents through these cells and the subsequent initiation of the heartbeat. However, as the human heart DT, mechanistic modeling requires massive and accurate simulation parameters, which are particularly challenging to acquire from biological entities in the human body, i.e., from molecular to organ level, and are typically limited to only a subset of all available biomolecular [105]. GAI techniques can overcome these challenges, through learning the underlying distribution and sequential or temporal relations of data for modeling. In the following, we provide a survey of the application of GAI in HDT digital modeling, including digital modeling of human cells, tissues and organs, as summarized in Table VI.
TABLE VI: Summary of GAI-based Approach for Digital Modeling in HDT.
@@ III-E GAI-enabled Data Analysis
Data analysis is the crucial component in HDT, which analyzes the data collected from the physical world and the data generated in the digital world. The results from data analysis are essential information in providing HDT services, such as diagnosis and prescription. Traditional data analysis methods, such as CNN-based [118] and U-Net-based methods [119], highly rely on a large amount of labeled data and difficult in handling high-dimensional HDT data. To this end, GAI has showcased remarkable potential in data analysis, including classification, segmentation, anomaly detection and prediction, as summarized in Table VII.
TABLE VII: Summary of GAI-based Approach for Data Analysis in HDT.
@@ IV-A Personalized Health Monitoring and Diagnosis
GAI has been successfully applied in anomaly detection from the collected patient data in HDT. For example, Wang et al. in [135] proposed a GAN-based approach for ECG signal analysis, achieving automatic cardiac diagnosis, as shown in Fig. 11. Specifically, the proposed approach involved two-level hierarchical deep learning framework with GAN. The first-level model was composed of a memory-augmented deep autoencoder with GAN (MadeGAN), which aimed to differentiate abnormal signals from normal ECGs for anomaly detection. The second-level learning aimed at robust multi-class classification for different arrhythmia identification, which was achieved by integrating the transfer learning technique to transfer knowledge from the first-level learning with the multi-branching architecture to handle the data-lacking and imbalanced data issues. Experimental results showed that the proposed approach can effectively capture the disease-altered feature patterns from ECG signals, yielding better performance in predicting heart disease compared to other existing methods. Moreover, this hierarchical deep learning framework can be broadly implemented to study other patient data, such as EEG and PPG, for smart anomaly detection. Following this, Another GAI technique, VAE, has been applied in this field. Staffini et al. in [136] proposed a disentangled VAE with a bidirectional LSTM (BiLSTM) backend to detect in an unsupervised manner anomalies in heart rate data collected during sleep time with a wearable device. Specifically, the added BiLSTM backend to the VAE model allowed it to capture contextual relationships in VAE-processed heart rate sequences by analyzing the information flow’s forward and backward directions. The proposed approach can better model the considered time series and learn more accurate patterns. Experimental results showed that the proposed approach outperformed five well-known algorithms in terms of anomaly detection from the heart rate data.
Beyond cardiac diseases, GAI has demonstrated successful applications in anomaly detection for monitoring and diagnosing other conditions like breast cancer [137] and Alzheimer’s [138]. These advancements utilize diverse patient data collected within HDT. Thus, GAI-driven HDT holds the promise of delivering precise, timely, and proactive personalized health monitoring and diagnosis, significantly enhancing patient care and overall well-being.
@@ IV-B Personalized Prescription
GAI has been successfully applied in treatment outcome predictions and simulations inside the HDT for personalized prescriptions. For instance, Jarada et al. in [139] proposed a VAE-based approach, called similarity network fusion-collective VAE (SNF-CVAE), for predicting drug-disease interactions, as shown in Fig. 12. Specifically, SNF-CVAE integrated similarity measures, similarity selection, SNF, and CVAE to conduct a non-linear analysis and improve the drug-disease interaction prediction accuracy. Additionally, to further demonstrate the reliability and robustness of SNF-CVAE, the authors conducted two case studies on the top predicted drug candidates for potentially treating Alzheimer’s disease and Juvenile rheumatoid arthritis, which were successfully validated against clinical trials and published studies. In addition, GAN has also been applied in this field. Xu et al. in [140] introduced a GAN-based method for predicting the short-term therapeutic outcomes of anti-vascular endothelial growth factor (VEGF) therapy in treating diabetic macular edema (DME). They employed the state-of-the-art GAN-style algorithm, pix2pixHD [141], renowned for progressively synthesizing high-resolution and realistic images. This approach was utilized to generate post-therapeutic optical coherence tomography (OCT) images. Their experimental results showcased that the established pix2pixHD model effectively produced post-therapeutic OCT images close to the ground truth, providing credible insights to aid ophthalmologists in predicting and assessing the response to anti-VEGF treatment.

### v2

@@ III-B GAI-enabled Communication
HDT relies on real-time data transmission to keep synchronization between any PT-VT pair to ensure high-fidelity of VT [82, 83, 84]. This synchronization is, however, data-driven and delay-sensitive. Furthermore, data acquired in the physical world is often massive and complex. In addition to real-time synchronization, interacting with VTs involves more complex information that needs to be transmitted between the PT and users in the physical world. Multi-modal information, such as 3D virtual items, text, images, haptic signals, among others, needs to be transmitted in HDT under various applications, such as the virtual surgery, to enhance the immersive experience. These specific characteristics place a significant burden on current communication networks [85]. To address these issues, this subsection delves into the GAI-aid semantic communication [67] and cross-modal communication [69] for communication in HDT, as summarized in Table V.
TABLE V: Summary of GAI-based Approach for Communication in HDT.
@@ III-C GAI-enabled Data Management
Data management is a crucial step in the successful implementation of HDT. First, the collected multi-source data in HDT may have the characteristics of heterogeneity, multi-scale and high noises. Hence, as the critical step in data management, data pre-processing is indispensable in HDT, such that issues like missing data and noise data can be properly handled, and thereby providing high-quality data for downstream tasks in HDT [13]. Moreover, the data in HDT are highly sensitive, especially for individual-level data, and any leakage may result in serious ethical and moral concerns. Therefore, effective security and privacy schemes are imperative to protect data in HDT. In the following, we will discuss the application of GAI in data pre-processing, followed by its application in security and privacy schemes, as summarized in Table VI.
TABLE VI: Summary of GAI-based Approach for Data Management in HDT.
@@
Traditional data imputation methods, such as K-nearest neighbors-based [95] and deep learning-based imputation [96], often rely on existing observed data for modeling and prediction, and cannot directly generate new data samples. This will make it hard for traditional data imputation methods to handle missing severe data issues, such as data modality missing. To this end, GAI, with its powerful generative capabilities, can effectively handle the missing data issues in HDT. For instance, with the potential of substituting missing data accurately and efficiently, Dong et al. in [97] used GAN to impute missing values in large clinical datasets collected from PTs with mixed-type variables. Specifically, the method adopted by authors was generative adversarial imputation nets (GAIN), where the generator observed some components of a real clinical data vector, imputed the missing components conditioned on what was actually observed, and outputs a complete clinical data vector. The discriminator then took a completed clinical data vector and attempted to determine which components were observed and which were imputed. Additionally, to ensure that discriminator forced generator to learn the desired data distribution, the authors provided discriminator with some additional information in the form of a hint vector. The accuracy of GAIN was measured by imputation error, defined as the difference between the imputed values and real value. Specifically, normalized root mean square error (NRMSE) was used to assess difference for continuous variables, while proportion of falsely classified (PFC) was used for categorical variables. The experimental results showed that GAIN outperformed the traditional imputation models, MICE [98] and missForest [99], in terms of NRMSE and PFC in the imputation of missing data in clinical datasets, under the missingness rate of 20% and 50%. Acquiring multi-modality data is crucial for the implementation of HDT. However, the missing modalities may happen in some conditions, such as failure in data transmission. Fortunately, diffusion models have shown favorable results for generating missing modalities utilizing cross-modalities and producing ones using other modality types. For instance, Lyu et al. in [100] proposed a GAI framework, called diffusion and score-matching models, which took advantage of the recently introduced denoising diffusion probabilistic models (DDPMs) [101] and score-based diffusion models [102], for translating MRI to CT. Specifically, they presented conditional DDPM and conditional stochastic differential equation (SDE) [103], where their reverse process was conditioned on T2w MRI images. The authors adopted the DDPM and SDE with three different sampling methods, namely Euler-Maruyama, Prediction-Corrector, and probability flow ordinary differential equation. In their experiments, they utilized the structural structural similarly index measure (SSIM) and peak signal-to-noise ratio (PSNR) metrics to evaluate their quality of translating MRI to CT. The datasets used in this experiment is Gold Atlas male pelvis dataset [104], and the results showed that the proposed method achieved SSIM = 0.86 and PSNR = 27.31, which was superior than both GAN (SSIM = 0.83, PSNR = 26.84) and CNN-based methods (SSIM = 0.84, PSNR = 26.98) [105]. Similarly, to cope with the missing modality issue, Meng et al. in [60] proposed a unified multi-modal conditional score-based generative approach (UMM-CSGM), which synthesized the missing modality based on all remaining modalities as conditions. The proposed model was a conditional SDE format, employing only a score-based network to learn different cross-modal conditional distributions. The experimental results showed that the UMM-CSGM could generate missing-modality images with higher fidelity and structural information of the brain tissue compared to GAN-based methods [106, 107, 108, 109, 110].
Noise data is a common issue during the data acquisition in HDT. Noise reduces the data quality and is especially significant when the point of interests are minor and have relatively low contrast, which hinders the downstream tasks in HDT [111]. Traditional data denoise methods, such as CNN-based [112] and U-Net-based methods [113], highly rely on a significant amount of labeled training data to build the denoising models. However, obtaining such data is challenging, especially at the individual level. To this end, with the strong generative abilities of GAI, it is convenient for diverse denoising problems in HDT data. For instance, PET is a medical imaging technique used to detect metabolic activities within the human body. The information provided by PET can be used to update or calibrate the VT models. However, due to various physical degradation factors, PET often suffers from low SNR and resolution. To denoise the PET images, a DDPM-based method has been proposed for PET image denoising. Gong et al. in [58] proposed the DDPM-based framework for PET denoising, called PET-DDPM, which collaborated with an assistive modality embedding as prior information to DDPM formulation. Quantitative results demonstrated that PET-DDPM with SSIM = 0.91 and PSNR = 33. 84 outperformed U-Net based denoising networks with SSIM = 0.87 and PSNR = 33.32 [114], showcasing its superior performance in PET denoising. For denoising MRI that with severely SNR-limited, Xiang et al. in [115] proposed a self-supervised denoising method, called denoising diffusion models for denoising diffusion MRI (DDM2). Specifically, their approach consisted of three stages, which integrated statistic-based denoising theory into diffusion models and performed denoising through conditional generation. During inference, they represented input noisy measurements as a sample from an intermediate posterior distribution within the diffusion Markov chain. The experiment results showed that DDM2 demonstrated superior denoising performances ascertained with clinically-relevant visual qualitative and quantitative metrics than Noise2Score (N2S) [116] and other denoising methods. Specifically, DDM2 not only suppressed noise the most but also restored anatomical details commented by two neuroradiologists. Additionally, compared to N2S, DDM2 demonstrated an average of 0.95 and 0.93 improvement on SNR and Contrast-to-Noise Ratio, respectively.
@@
HDT environments may suffer from various security threats, such as eavesdropping attacks in data transmission between PTs and VTs causing privacy leakage. A typical privacy protection mechanism for HDT data is anonymization [117, 118]. However, this method can easily compromise the data distribution, and it is hard to balance data privacy and usability. To this end, GAI can generate synthesized HDT data, which maximally preserves original data distribution while guaranteeing data security and privacy. For instance, to lower the risk of breaching individual confidentiality during data sharing in HDT, Yoon et al. in [119] proposed a GAN-based framework, called anonymization through data synthesis using GAN (ADS-GAN), for generating synthetic data that closely approximates the distribution of variables in an original EHR dataset, achieving anonymization of EHR, as shown in Fig. 6. Specifically, ADS-GAN generated the synthetic data conditioned on the original data, and different from traditional conditional GAN framework, conditioning variables in ADS-GAN were not pre-determined but instead were optimized from real EHR. Additionally, authors proposed a mathematical definition for identifiability to quantify the degree of anonymization. Particularly, $\epsilon$-identifiability they defined meant that the probability of patients can not identifiable by the synthetic patient data was larger than $\geq 1-\epsilon$, which was attributed to the synthetic patient data were “different enough” from the real ones. To prevent re-identification by generated synthetic data, authors integrated the constraint, $\epsilon$-identifiability, into ADS-GAN. To this end, they designed a identifiability loss for training discriminator in ADS-GAN. The experiment results showed that ADS-GAN performed superior than traditional conditional GAN frameworks, i.e., PATE-GAN [120] and WGAN-GP [105]. ADS-GAN can be 27% and 67% better than PATE-GAN in terms of both Jensen-Shannon Divergence (JSD) and the Wasserstein distance with the same identifiability. While with the same quality of the generated samples, ADS-GAN reduced the identifiability by 41% compared to WGAN-GP.
@@ III-D GAI-enabled Digital Modeling
The classic HDT digital modeling technology, mechanistic modeling, has pioneered the integration of biology and physiology domain knowledge to allow robust and accurate modeling [121]. For instance, a human heart DT model has been developed, which consisting of approximately 100 million virtual heart cell patches, with each patch modeled by around 50 equations [122]. The human heart DT accurately represented the interconnected cardiac muscle cells, effectively simulating the transmission of electrical currents through these cells and the subsequent initiation of the heartbeat. However, as the human heart DT, mechanistic modeling requires massive and accurate simulation parameters, which are particularly challenging to acquire from biological entities in the human body, i.e., from molecular to organ level, and are typically limited to only a subset of all available biomolecular [121]. GAI techniques can overcome these challenges, through learning the underlying distribution and sequential or temporal relations of data for modeling. In the following, we provide a survey of the application of GAI in HDT digital modeling, including digital modeling of human cells, tissues and organs, as summarized in Table VII.
TABLE VII: Summary of GAI-based Approach for Digital Modeling in HDT.
@@ III-E GAI-enabled Data Analysis
Data analysis is the crucial component in HDT, which analyzes the data collected from the physical world and the data generated in the digital world. The results from data analysis are essential information in providing HDT services, such as diagnosis and prescription. Traditional data analysis methods, such as CNN-based [134] and U-Net based methods [135], highly rely on a large amount of labeled data and difficult in handling high-dimensional HDT data. To this end, GAI has showcased remarkable potential in data analysis, including classification, segmentation, anomaly detection and prediction, as summarized in Table VIII.
TABLE VIII: Summary of GAI-based Approach for Data Analysis in HDT.
@@
III-F Promoting the GAI-driven HDT
In order to successfully integrate GAI-driven HDT technology into IoT-healthcare, it is vital to involve both patients and healthcare professionals in the process. Here we provide a few strategies for engaging them in the development and adoption of GAI-driven HDT.
1) Educational workshops and training: Start by building knowledge about GAI-driven HDT among healthcare professionals. Educating them on the benefits and limitations, explaining how it can be effectively used, discussing privacy concerns, and any possible adverse effects, can enable healthcare professionals to use it more effectively and also communicate these aspects convincingly to the patients.
2) Collaborative development: Involve healthcare professionals in the development process of these applications to ensure they align with current workflow and improve healthcare delivery. In turn, the developer should also do navigation for understanding the demands and concerns of patients. Their first-hand insights are invaluable to make the technology even more helpful and user-friendly.
3) Ethical and privacy concerns: To mitigate the potential ethical and privacy issues, there needs to be a solid foundation of transparency. Ethical AI is imperative to be considered. There needs to be a solid foundation of transparency. The GAI-driven HDT technology providers should be able to provide clear explanations of its processes and decisions to both patients healthcare professionals. This means explicit consent should be obtained before any personal health data is collected or used. The GAI-driven HDT should be trained and tested to ensure that it does not exhibit any bias, ensuring fairness and equity in healthcare provisioning. Additionally, accountability and oversight are needed to ensure the individuals or the entities responsible for the GAI-driven HDT making its corresponding decisions accountable. Besides, the strategies for ensuring data privacy and security are crucial. Technologies such as blockchain [151], federated learning [152], differential privacy [153], secure multi-party computation [154], and physical layer security aware wireless communications [155, 156, 157], can be designed and implemented for GAI-driven HDT to prevent unauthorized access and data breaches.
@@ IV-A Personalized Health Monitoring and Diagnosis
GAI has been successfully applied in anomaly detection from the collected patient data in HDT. For example, Wang et al. in [158] proposed a GAN-based approach for ECG signal analysis, achieving automatic cardiac diagnosis, as shown in Fig. 11. Specifically, the proposed approach involved two-level hierarchical deep learning framework with GAN. The first-level model was composed of a memory-augmented deep autoencoder with GAN (MadeGAN), which aimed to differentiate abnormal signals from normal ECGs for anomaly detection. The second-level learning aimed at robust multi-class classification for different arrhythmia identification, which was achieved by integrating the transfer learning technique to transfer knowledge from the first-level learning with the multi-branching architecture to handle the data-lacking and imbalanced data issues. The performance of the GAN-based anomaly detection was evaluated by two global performance metrics, i.e., area under receiver-operating-characteristic curve (AUROC), and area under precision-recall curve (AUPRC). The ROC characterizes the relationship between false positive rate (FPR) and true positive rate (TPR). Then, AUROC is a score quantifying the general prediction performance across all thresholds. Meanwhile, the AUPRC evaluates the general relationship between recall and precision. In the GAN-based anomaly detection, the performance of the proposed MadeGAN was superior than traditional methods, i.e., AutoEncoder (AE) and BeatGAN [159]. Specifically, MadeGAN achieved the best performance with AUROC and AUPRC scores of 0.954 and 0.936. In addition, MadeGAN improved AUROC by 5.3% compared with AE (0.906), and by 1.4% compared with BeatGAN (0.941). Furthermore, the improvement of AUPRC is more significant, with an increase of 5.6% and 1.5% compared with AE (0.886) and BeatGAN (0.922), respectively.
In addition, the industry is also walking on this way, aiming to facilitate the implementation of GAI-driven HDT. The global leader in GAI for pathology, Paige, leverages large and diverse dataset in cancer diagnostics to develop their GAI-based products for cancer pathology detection and analysis [160]. Their GAI-based products has been independently validated and clinically utilized by hospitals and labs around the world, and the reduction in cancer detection errors reaches 70% [161], while the reduction in time to diagnosis is 65.5% [162]. It is expected that such powerful tools can be integrated into HDT for personalized health monitoring and diagnosis.
Thus, GAI-driven HDT holds the promise of delivering precise, timely, and proactive personalized health monitoring and diagnosis, enhancing patient care and overall well-being.
@@ IV-B Personalized Prescription
GAI has been successfully applied in treatment outcome predictions and simulations inside the HDT for personalized prescriptions. For instance, Jarada et al. in [163] proposed a VAE-based approach, called similarity network fusion-collective VAE (SNF-CVAE), for predicting drug-disease interactions, as shown in Fig. 12. Specifically, SNF-CVAE integrated similarity measures, similarity selection, SNF, and CVAE to conduct a non-linear analysis and improve the drug-disease interaction prediction accuracy. For evaluating the performance of SNF-CVAE in drug-disease interactions, authors introduced five evaluation metrics, i.e., precision, recall, F1-score, area under curve-receiver operator characteristic (AUC-ROC), and area under curve-precision-recall (AUC-PR) [164]. Specifically, Precision $=TP/(TP+FP)$, Recall $=TP/(TP+FN)$, and F1-score $=2*(Precision*Recall)/(Precision+Recall)$. $TP$ was predicted and known drug-disease interactions, $TN$ was unpredicted and unknown drug–disease interactions, $FP$ was predicted and unknown drug–disease interactions, and $FN$ was unpredicted and known drug–disease interactions. Additionally, AUC-ROC can be described as plotting the $TP$ against $FP$, while AUC-PR can be described as a trade-off between Precision and Recall. The simulation results showed that SNF-CVAE achieved significantly higher accuracy with Precision = 0.902, Recall = 0.883, F1-score = 0.893, AUC-ROC = 0.958 and AUC-PR = 0.970, compared the traditional methods, Pareto dominance and collaborative filtering-based drug-disease prediction [165] with Precision = 0.487, Recall = 0.085, F1-score = 0.144, AUC-ROC = 0.542 and AUC-PR = 0.225.
Beyond the academic successes, the industry has also achieved some milestones in this area. The pioneer in utilizing GAI for drug discovery and development, Insilico Medicine, had developed a GAI powered commercially available platform, Pharma.AI [166]. Their lead drug, INS018_055, was discovered and designed using Pharma.AI, which was designed to treat the devastating lung condition idiopathic pulmonary fibrosis (IPF) [167]. IPF is a disease marked by excessive scarring of the lungs that affects 5 million people worldwide and has few treatment options. INS018_055 was nominated as a preclinical candidate – a record 18 months from target discovery — and quickly accelerated into further trials, from first-in-human, to Phase I trials with healthy volunteers, where the drug showed excellent results, delivering a favorable pharmacokinetic and safety profile. Now (i.e., March 2024), the drug is currently in Phase 2a trials in both the U.S. and China with IPF patients – less than 4 years from discovery. Thus, compared to traditional drug discovery methods, which typically takes from 10 to 15 years to develop a drug, the GAI-driven methods presented here has significantly shorten the process. Promisingly, Pharma.AI can be integrated into the development of GAI-driven HDT, which can acts the human digital testbed by simulating the drug-disease interactions.